OpenStudy (anonymous):

Express 9 ×9 × 9 × 9 × 9 in exponential form.

jimthompson5910 (jim_thompson5910):

how many copies of '9' are there

OpenStudy (anonymous):

5 soo im guessing 9^5

jimthompson5910 (jim_thompson5910):

Exactly, \[\Large 9 \times 9 \times 9 \times 9 \times 9 = 9^{5}\]

jimthompson5910 (jim_thompson5910):

exponents are shortcuts for repeated multiplication (in much the same way multiplication is a shortcut for repeated addition)

OpenStudy (anonymous):

okay (:

OpenStudy (anonymous):

9^5